OpenStudy (anonymous):

P 8 I think p=4 - = - All answers p=2 3 12 p=8 p=12 p=24

OpenStudy (alexandervonhumboldt2):

p/8=3/12

OpenStudy (alexandervonhumboldt2):

p=1/4*8 p=2

OpenStudy (alexandervonhumboldt2):

p=2

OpenStudy (anonymous):

Man I get so close. each time Thanks